OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11

hero (hero):

y = (3-3x)/-6 y = (11+7x)/5 y = y -6(11+7x) = 5(3-3x) -66 -42x = 15 - 15x 66+15 = -42x + 15x 81 = 27x 81/27 = x 3 = x

OpenStudy (anonymous):

x-3

OpenStudy (radar):

I believe a check will reveal that x=-3. Lets check. If x=-3, then 3(-3)-6y=3 or -9-6y=3. Solve for y. -6y=3+9 or -6y=12 then y=-2. Using the values x=-3, and y=-2 substitute these values into the original equations. 3(-3)-6(-2)=3 -9+12=3 3=3 Check! 7(-3)-5(-2)=-11 -21+10=-11 -11=-11 Again a Check! There is an error in somewhere, and I think it is a sign error.